OpenStudy (anonymous):

7 years ago, Samir was 3 times as old as Deepak. In 4 years, Samir will be twice as old as Deepak. What is Deepak’s present age?

OpenStudy (anonymous):

are there any answer options?

OpenStudy (anonymous):

Deepak is 18 years old @yomamabf

OpenStudy (anonymous):

can i get a medal?

OpenStudy (anonymous):

uh no cuz u googled it and didnt even give the steps

OpenStudy (anonymous):

you want me to expain my answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Present Deepak: D + 7 Samir: 3D + 7 4 years from now D's age: D+11 S's age: 2(D+11) Now we just need to represent this with an equation. What is one more piece of information that we know that isn't represented by this table? Well, we know that people age at the same rate. So Samir's age 4 years from now will be 11 more than his age was 7 years ago. So we can write: 3D + 11 = 2(D + 11) And then we can solve for D: 3D + 11 = 2(D + 11) 3D + 11 = 2D + 22 D + 11 = 22 D = 11 Then, remember that D = Deepak's age 7 years ago. So that means 11 + 7 = 18, his age now.

OpenStudy (anonymous):

u copied and pasted this from the magoosh website

OpenStudy (anonymous):

magoosh?

OpenStudy (anonymous):

give me medal